0 2y+3 S4) ss x- 2y X+ y ୧ dxdy 1-y (x-2y)² integralinin değerini bulunuz. (15 puan)
Answer #1

Solm. je 27 +3 aty (x-24) ex-2y da dy - y y=-1 to y=0 and a=-y x=-y to x= 2y+3 = x+y= to x- 2y=3 let x+y = U, x-2y = - y=u-ce